Graham, Liz and Mel met to check on stores and general status.  Conditions overcast and cold – 7 degrees.  Previous day had been warm enough for bees to emerge for cleansing flights.
Hive #2 – Fondant added. Super removed.
Hive #3 – Feeding not required.
Hive #4 – Small amount of fondant added.
Hive #5 – High number of dead bees behind mouseguard but general status OK. Fondant added.
Hive #6 – Fondant added. Empty super removed.
Hive #7 – Plenty of stores remaining so no need to feed.  Low number of bees.
Hive # 8 – Fondant added but stores still available in super.
Hive # 15F – Plenty of stores still available.  Colony had been bad-tempered last season so consider moving out of shade to more open position in apiary.
Hive #16B – Fondant added although stores still available.
Next custodians meeting Sunday 26 Feb at 1:00pm.